Two cards drawn from a standard deck of cards . Find the probability of,
P(ace,then king) if replacement occurs
P(heart,then club) if no replacement occurs

Answer by edjones(8007)

P(ace,then king) if replacement occurs:
4/52 * 4/52 = 1/169
.
P(ace,then king) if no replacement occurs:
4/52 * 4/51 = 4/663
.
Ed